You are recently hired as a staff accountant for a small finished goods manufacturing company. Part of your duties include doing the month end inventory of finished goods. After a few months you do not look forward to this as the amount of inventory seems to be increasing. In order to satisfy your thoughts on this increase of inventory you decide to review the financial information for the last few months.

Looking over the Income Statement you see the profits have been steady, but the gross profit percentage has increased, and the cost of goods sold have decreased. This does not seem possible as the company has increased the amount of inventories.

Identify why this situation could exist, providing an explanation which can be given to the CFO.

In order to assist in controlling the costs and providing a lower inventory carrying cost, select a costing system and explain why it should be utilized.
Answer #1

Lets start by understanding this amount from the begining. sales - COGS = Gross profit.

Now, it is evident from the above formula that when the cost of goods sold decreases, the amount of gross profit would increase.

In the above case, profits have been steady, but the gross profit percentage has increased, and the cost of goods sold have decreased. This does not seem possible as the company has increased the amount of inventories. The company might be using First in First out method for inventory costing. for example, company bought four materials A,B,C,D, in the same order. the cost was A 10, B 6, C 4, D 4, and sales price was 20. As per FIFO method, if A was bought first, it would be assumed that A was sold first. that means, initially, thee cost of inventory would be high, But for B C D, the cost would be lower and thus this would result in increase in Gross Profit as compared to Gross profit in case of material A. Gross Profit in case of A is 20-10= 10, B 20-6=14, C 20-4=16, D 20-4=16.

IAS 2 Inventories contains the requirements on how to account for most types of inventory. The standard requires inventories to be measured at the lower of cost and net realisable value (NRV)